How many coconuts did the first two monkeys take?


1
Expert's answer
2021-10-12T09:18:38-0400

Have many coconuts are there in two heaps when the second money goes home?


"12-2=10"

Have many coconuts are there when  the second monkey awakes?


"10(\dfrac{3}{2})+1=16"


Have many coconuts are there when  the first monkey awakes?


"16(\dfrac{3}{2})+1=25"


Have many coconuts did the first monkey take?


"(25-1)(\dfrac{1}{3})=8"

Have many coconuts did the second monkey take?


"(16-1)(\dfrac{1}{3})=5"

How many coconuts did the first two monkeys take?


"8+5=13"

The first two monkeys took 13 coconuts.